OpenStudy (pphalke):

In a​ triangle, the measure of the first angle is three times the measure of the second angle. the measure of the third angle is 80 degrees more than the measure of the second angle. use the fact that the sum of the measure of the three angles of a triangle is 180 to find the measure of each angle

OpenStudy (dayakar):

let the second angle =x first angle = 3x third angle = 80 sum of the three angles in a triangle = 180 can u find the angles?

OpenStudy (anonymous):

@dayakar Third angle is 80+x

OpenStudy (helasi95):

wouldnt it be: 3x + x + (x+80)= 180 ?

OpenStudy (dayakar):



OpenStudy (dayakar):

3x+x+80 =180 4x +80=180 4x= 180-80 4x= 100 what is x ?

OpenStudy (anonymous):

@Helasi95 Yes 3x + x + (x+80)= 180 is right. @dayakar Is forgetting that.

OpenStudy (dayakar):

i am sorry @Zelda is correct

OpenStudy (helasi95):

3x + x+ (x+80) = 180 5x + 80 =180 5x = 100 x=100/5 x=20 ...Then: 1. 3x = 60 2. x = 20 3. x + 80 = 100 100 + 60 + 20 = 180.

OpenStudy (dayakar):

3x+x +80+x =180 5x +80 =180 5x =180-80 5x =10 what is x ?

OpenStudy (helasi95):

Did you mean: 5x = 100, what is x? Well... the answer is up x = 100/5 = 5 x 10= 50 + (5 x 10) = 20 x = 20.